the sum of two numbers is 22 and the sum of their square is 250. find the numbers

Answer by Alan3354(69443)

the sum of two numbers is 22 and the sum of their square is 250. find the numbers
---------
x + y = 22
x^2 + y^2 = 250
----------
y = 22-x
Sub for y
x^2 + (22-x)^2 = 250
x = 9
y = 13
=============